OpenStudy (anonymous):

The diameter of a soup can is about 7cm. The label wraps around the can. About how many cm is the label to go all the way around the soup can? Use 22/7 for (pie). Explain how you are reasonable.

OpenStudy (anonymous):

They are asking, basically, for the circumference. Well C=pi(d) C=(22/7)(7)=22, so you would need 22 cm

OpenStudy (anonymous):

Thank you very much!